OpenStudy (anonymous):

What is an equation of a parabola with the given vertex and focus? vertex: (–2, 5); focus: (–2, 6)

OpenStudy (anonymous):



OpenStudy (anonymous):

since both the \(x\) values are the same, you know the \(x\) term is squared also because the focus is above the vertex, you know if faces up see the lousy picture

OpenStudy (anonymous):

equation therefore has the form \[(x-h)^2=4p(y-k)\] where \((h,k)\)is the vertex so you know it looks like \[(x+2)^2=4p(y-5)\] and all you need is \(p\)

OpenStudy (anonymous):

from 5 to 6 is one unit up, making \(p=1\) and so \(4p=4\)

OpenStudy (anonymous):

and your equation is \((x+2)^2=4(y-5)\) you can check it here http://www.wolframalpha.com/input/?i=+parabola+%28x%2B2%29^2%3D4%28y-5%29

OpenStudy (anonymous):

Sorry I had to step away from the computer. I'm still a bit confused. What's the outcome of this is supposed to look like?